OpenStudy (anonymous):

A tanker used to transport oil is filled up by two pipes in 4 hours.If the larger pipe alone takes 6 hours less than the smaller pipe to fill up the tanker, find the time in which the tanker will be filled by each pipe operating individually.

OpenStudy (anonymous):

6 hrs

OpenStudy (anonymous):

can you explain how you did it?thanks a lot man.

OpenStudy (anonymous):

(1/x)+(1/x+6)=1/4

OpenStudy (anonymous):

(1/x)+[1/(x+6)]=1/4

OpenStudy (anonymous):

I don't get it.Can you put it in a simpler way??

OpenStudy (anonymous):

Let , larger pipe takes x hrs , individually. so ,smaller pipe takes x+6 hrs, Formula , if both pipes are operating at same time, total time takes = (1/x)+[1/(x+6)]=1/4

OpenStudy (anonymous):

why do we need 1/x and not just x? sorry i am quite weak at math

OpenStudy (anonymous):

1/x is taken because volume of the oil is considered 1. you can take any variable too

OpenStudy (anonymous):

got it yet legend ?

OpenStudy (anonymous):

let the larger pipe takes x hr and smaller takes x+6 hrs the volume be v in 4 hrs the larger pipe fills 4v/x and smaller 4v/(x+6) so 4v/x+4v/(x+6)=v 1/x+1/(x+6)=1/4 4(2x+6)=x^2+6x x^2-2x-24=0 (x-6)(x+4)=0 so x=6 so the larger one fills in 6 hrs and the smaller in 6+6=12 hrs

OpenStudy (anonymous):

yeah got it..Thanks dispankarstudy.

OpenStudy (anonymous):

6h and 12 is correct answer

OpenStudy (anonymous):

yupp